# Asymmetric Contrastive Multimodal Learning for Advancing Chemical Understanding

### v3

@@ 4.2 Graph encoder
Graph pooling and readout. The node representations at the final iteration will be processed to generate the fixed-length graph-level representation $\mathbf{h}_{g}$, which involves graph pooling and readout operations. The pooling layer is usually applied to get a coarser graph which can be further reduced within a readout function, which is similar to conventional CNNs [68]. There is no strict distinction between pooling and readout operations in GNNs. Despite some studies [69, 70] that leveraged stacks of pooling layers for coarse graphs, most widely used works handling pooling and readout simultaneously by applying a permutation invariant function directly on all nodes in the graph to generate a fixed length graph representation [71, 72, 73, 74], which is generally formalized as
@@
The choice of aggregation operator. Many previous works [67, 51, 73, 75] demonstrate that the choice of aggregation function contributes significantly to the expressive power and performance of the model. In this work, we used sum operator in both message aggregation and graph readout function, which enables simple and effective learning of structural graph properties implied in [51].
4.3 Multimodal contrastive learning
@@
5 Experimental details
5.1 Datasets
5.1.1 Training Datasets
For the molecular depiction image dataset, the molecules come from the Natural Products Magnetic Resonance Database (NPMRD) [79] and the images were produced by RDKit [80]. In total, there are about 270,000 samples in this dataset. For the SMILES dataset, molecules and SMILES come from NPMRD as well. In total, there are about 270,000 samples in this dataset. For 1H NMR dataset and 13C NMR dataset, the molecules, 1H NMR spectra, and 13C NMR spectra come from NPMRD as well. There are about 18,000 samples in 1H NMR dataset and 18,000 samples in 13C NMR dataset. For the GCMS dataset and LCMS dataset, molecules, LCMS spectra, and GCMS spectra come from MassBank of North America (MoNA) [81]. There are about 10,000 samples in the GCMS dataset and 13,000 samples in the LCMS dataset.
5.1.2 Zero-Shot Datasets
@@
For the molecular candidate pool, 1M molecules were randomly chosen from PubChem, no intersection with the training dataset. For the spectra, 1000 spectra were used from the validation dataset.
5.2 Training Details
5.2.1 Pretraining in ACML
@@
For graph modality, each atom is embedded by its atomic number and chirality type, and each bond is embedded by its bond type and stereotype. We used GNNs with ReLU activation function and a sum pooling is applied on each graph as the final readout operation to extract the hidden graph representation. The best model settings are provided in Table 4. Besides GIN [51], we tried multiple convolution techniques and found that GCN [65] achieved similar results to GIN, whereas GAT [66] and GraphSage [67] degraded performance during validation. We noticed that the number of layers did not have a strong effect as long as it exceeded a value of 5, implying that a shallow graph encoder is expressive enough. For chemical modalities, we used pre-trained unimodal encoders according to Table 3, where each encoder keeps frozen in the training phase. Outputs from the graph encoder and the chemical unimodal encoder are projected into the same dimensional joint hidden space using a 2-layer MLP with a specific projection dimension as provided in Table 4. All models were optimized with AdamW [82] in Pytorch implementation with $\beta_{1}=0.9,\beta_{2}=0.999$, and weight decay of 0.001. The model is trained with a batch size of 128 for a total of 100 epochs. Each model is trained on one NVIDIA Tesla V100 GPU using float32 precision. The entire framework is training efficiently since only one shallow graph encoder and two MLPs from the projection module need to be updated through training.
@@
G-1H NMR GIN 3 512 256
G-13C NMR GIN 3 512 128
G-GCMS GIN 5 128 128
G-LCMS GIN 5 128 128
Table 4: Best Model settings in ACML.
5.2.2 Finetuning on molecular property prediction tasks
The following configurations were applied to all training tasks on the seven molecular benchmarks. We used a batch size of 32 and a maximal training epoch of 100. All experiments were conducted on one Tesla V100 GPU. We used the Adam optimizer. We tried different values of learning rate within [1e-2, 1e-3, 1e-4, 1e-5] and reported the best results. Before training, we performed data cleaning to remove certain molecules that failed to pass the sanitizing process in the RDKit or contained abnormal valence of a certain atom, as suggested in [83, 84, 44]. The detailed dataset statistics are summarized in Table 5.
@@ 5.2.2 Finetuning on molecular property prediction tasks
6 Competing interests
@@
7 Declaration of generative AI and AI-assisted technologies in the writing process
@@
8 Data availability
The pre-training data from NPMRD and MoNA, and the zero-shot PubChem dataset are publicly available online, well-organized, and can be directly downloaded. RDKit can assist in parsing the dataset into individual samples.
9 Code availability
The code accompanying this work is available Github: https://github.com/GainGod-Xu/ACMLProject.

### v4

@@ A.2 Graph Encoder
Graph Pooling and Readout.
The node representations at the final iteration will be processed to generate the fixed-length graph-level representation $\mathbf{h}_{g}$, which involves graph pooling and readout operations. The pooling layer is usually applied to get a coarser graph which can be further reduced within a readout function, which is similar to conventional CNNs [70]. There is no strict distinction between pooling and readout operations in GNNs. Despite some studies [71, 72] that leveraged stacks of pooling layers for coarse graphs, most widely used works handling pooling and readout simultaneously by applying a permutation invariant function directly on all nodes in the graph to generate a fixed length graph representation [73, 74, 75, 76], which is generally formalized as
@@
The Choice of Aggregation Operator.
Many previous works [69, 57, 75, 77] demonstrate that the choice of aggregation function contributes significantly to the expressive power and performance of the model. In this work, we used sum operator in both message aggregation and graph readout function, which enables simple and effective learning of structural graph properties implied in [57].
A.3 Multimodal Contrastive Learning
@@
Appendix B Experimental Details
B.1 Datasets
B.1.1 Datasets for Training ACML framework
For the molecular depiction image dataset, the molecules were sourced from the Natural Products Magnetic Resonance Database (NP-MRD) [81] and the images were produced by RDKit [82]. Similarly, the SMILES dataset, consisting of molecules and their corresponding SMILES representations, was also sourced from NP-MRD. In total, the dataset contains approximately 270,000 samples. For 1H NMR dataset and 13C NMR dataset, the molecules, 1H NMR spectra, and 13C NMR spectra were obtained from NP-MRD as well. There are about 18,000 samples in 1H NMR dataset and 18,000 samples in 13C NMR dataset. These NMR data are represented as 1D sequential data, consisting of peak locations and intensities. In detail, the raw NMR data contains the chemical shift locations and peak intensities. For 1H NMR, the chemical shift ranges from 0 - 10 ppm and for 13C NMR, the chemical shift ranges from 0 - 220 ppm. During 1D NMR processing, 1H NMR data is mapped onto a grid spanning 0 to 10 ppm with increments of 0.01 ppm, and 13C NMR data is mapped onto a grid spanning 0 to 220 ppm with increments of 0.1 ppm. The intensity corresponding to each chemical shift is recorded on the grid, with all other grid points set to 0. For the GCMS dataset and LCMS dataset, molecules, LCMS spectra, and GCMS spectra were sourced from MassBank of North America (MoNA) [83]. There are about 10,000 samples in the GCMS dataset and 13,000 samples in the LCMS dataset. We removed molecules with a graph size of 100 or more atoms, excluding hydrogen atoms from the graph size count. All data was accessed on July 15th, 2023, with no evidence of potential bias. To prevent data leakage in the validation dataset, duplicated spectra and molecules were removed.
B.1.2 Zero-Shot Datasets
@@
For the molecular candidate pool, 1M molecules were randomly chosen from PubChem [84] with access on July 15th 2023, no intersection with the training dataset. For the spectra, 1000 spectra were randomly chosen from the validation dataset. The data was accessed on July 15th, 2023, with no potential bias identified. To prevent data leakage, duplicated spectra and molecules were removed, and all data cleaning steps are fully documented.
B.2 Training Details
B.2.1 Pretraining in ACML
@@
For graph modality, each atom is embedded by its atomic number and chirality type, and each bond is embedded by its bond type and stereotype. We used GNNs with ReLU activation function and a sum pooling is applied on each graph as the final readout operation to extract the hidden graph representation. The best model settings are provided in Table 7. Besides GIN [57], we tried multiple convolution techniques and found that GCN [67] achieved similar results to GIN, whereas GAT [68] and GraphSage [69] degraded performance during validation. We noticed that the number of layers did not have a strong effect as long as it exceeded a value of 5, implying that a shallow graph encoder is expressive enough. For chemical modalities, we used pre-trained unimodal encoders according to Table 6, where each encoder keeps frozen in the training phase. Outputs from the graph encoder and the chemical unimodal encoder are projected into the same dimensional joint hidden space using a 2-layer or 3-layer MLP with a specific projection dimension as provided in Table 7. All models were optimized with AdamW [85] in Pytorch implementation with $\beta_{1}=0.9,\beta_{2}=0.999$, and weight decay of 0.001. The model is trained with a batch size of 128 for a total of 100 epochs. Each model is trained on one NVIDIA Tesla V100 GPU using float32 precision. The entire framework is training efficiently since only one shallow graph encoder and two MLPs from the projection module need to be updated through training.
Table 7: Best model settings in ACML.
@@
G-1H NMR GIN 5 512 256
G-13C NMR GIN 3 512 256
G-GCMS GIN 5 128 [128, 128]
G-LCMS GIN 5 128 [128, 128]
G-GCMS and G-LCMS use the 3-layer MLP in the projection module while others use the 2-layer MLP in projection.
B.2.2 Finetuning on Molecular Property Prediction Tasks
The following configurations were applied to all finetuning tasks from MoleculeNet [54] and TDC [55] accessed on July 15, 2023. For each task, we utilized a 5-layer GIN [57] as the backbone graph encoder, exploring hidden dimensions of [64, 128, 300]. The GIN encoder was initialized with weights pre-trained using the ACML framework. We used a batch size of 32 and a maximal training epoch of 100. We use scaffold splitting [86] to divide the dataset into the training set, validation set, and test set with a ratio of 8:1:1. Scaffold splitting helps avoid data leakage in molecular datasets because it ensures that structurally similar molecules do not appear in both the training and test sets. All experiments were conducted on one Tesla V100 GPU. We used the Adam optimizer. Different learning rates from [1e-2, 1e-3, 1e-4, 1e-5] are explored and the best results are reported. Before training, we performed data cleaning to remove certain molecules that failed to pass the sanitizing process in the RDKit or contained abnormal valence of a certain atom, as suggested in [87, 88, 44]. The detailed dataset statistics are summarized in Table 8.
@@ B.2.2 Finetuning on Molecular Property Prediction Tasks
B.2.3 Discussion of Training Efficiency
To validate our claim regarding training efficiency, we focused on the two phases: pretraining and fine-tuning.
Finetuning.
This stage is standardized across models, so the efficiency gains result solely from the graph architecture. We observed that in some datasets, a pretrained graph encoder with a smaller hidden dimension is already enough to achieve a good performance, as implied in Table 9. We believe this is because the proposed approach can learn good chemical knowledge through ACML pretraining.
Table 9: Embedding Dimensions of the Best Models for Different Datasets.
Embedding Dimension Datasets
64
BACE, BBBP, SIDER, ESOL, FreeSolv, HIA, PGP Inhibition,
CYP2C19 Inhibition, CYP2D6 Inhibition, CYP3A4 Inhibition
CYP1A2 Inhibition, CYP2C9 Inhibition, CYP2C9 Substrate
128
ClinTox, Tox21, HIV, CYP2D6 Substrate, CYP3A4 Substrate
300 PAMPA Permeability, Bioavailability
Pretraining.
We found the challenge of fairly comparing efficiency solely by time, given the different types of resource utilization involved, such as human efforts, dataset size, and CPU/GPU resources. Overall, the proposed approach uses a relatively small number of learnable parameters (only one graph encoder and two MLPs for projection), which is lightweight with minimal preprocessing and limited human involvement. For instance, SMILES-to-graph conversion takes just five minutes per million molecules on a standard laptop. In contrast, some baseline methods demand significant CPU or human resources for preprocessing (e.g., constructing motif vocabulary) or for multimodal pretraining, which requires generating augmented graphs and optimizing augmentation settings. Besides preprocessing efficiency, we provide a clear resource-efficiency comparison of pretraining. For MoMu and MotifConv, we report resource utilization as stated in their original papers. For L2P-GNN, MGSSL, MolCLR and the proposed ACML instantiations, we reproduced pretraining on a single NVIDIA Tesla V100 32GB GPU with 100 epochs. We reported both the dataset sizes and the training time per epoch in Table 10. These benchmarks indicate that our approach requires fewer resources, which is advantageous in scenarios where computational and financial costs are critical factors.
Table 10: Resource Utilization and Time per Epoch Comparison
Model GPU Requirement Dataset Size Time per Epoch
MoMu 8 V100, 32G $\sim$ 16K
MotifConv 8 RTX 2080, 11G Several hours per dataset
L2P-GNN 1 V100, 32G $\sim$ 250K $\sim$ 7 min
MGSSL 1 V100, 32G $\sim$ 250K $\sim$ 40 min
MolCLR 1 V100, 32G $\sim$ 1 Million $\sim$ 80 min
G-Image 1 V100, 32G $\sim$ 250K $\sim$ 3 min
G-SMILES 1 V100, 32G $\sim$ 250K $\sim$ 7 min
G-13C NMR 1 V100, 32G $\sim$ 18K $\sim$ 12 sec
G-1H NMR 1 V100, 32G $\sim$ 18K $\sim$ 12 sec
G-GCMS 1 V100, 32G $\sim$ 13K $\sim$ 8 sec
G-LCMS 1 V100, 32G $\sim$ 10K $\sim$ 7 sec
Appendix C Competing Interests
@@
Appendix D Declaration of Generative AI and AI-assisted Technologies in The Writing Process
@@
Appendix E Data Availability
The pre-training data from NP-MRD and MoNA, and the zero-shot PubChem dataset are publicly available online, well-organized, and can be directly downloaded. RDKit can assist in parsing the dataset into individual samples.
Appendix F Code Availability
The code accompanying this work is available Github: https://github.com/GainGod-Xu/ACMLProject. The supplementary materials S1 and S2 will be available when this work is formal published.